OpenStudy (anonymous):

Find the slope and y-intercept of the line. 14x + 4y = 24

sam (.sam.):

14x+4y=24 4y=-14x+24 y=-(7x-12)/(2) y=-7x/2 -6 Slope=-7/2, y-int=-6

OpenStudy (savvy):

slope = -7/2 y int. = 6

sam (.sam.):

y=-7x/2 +6 y-int=6